How to do problems in dipole moment in chemical bonding

Grade:11

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

Please find the link below

https://www.askiitians.com/iit-study-material/iit-jee-physics/chemical-bonding/covalency/#dipole-moment

Also ask a specific doubt where you are not able to solve the problem

Regards
TEJA KRISHNA
4 years ago
 
  • Dipole Moment

Difference in polarities of bonds is expressed on a numerical scale. The polarity of a molecule is indicated in terms of dipole moment (μ). To measure dipole moment, a sample of the substance is placed between two electrically charged plates. Polar molecules orient themselves in the electric field causing the measured voltage between the plates to change.

The dipole moment is defined as the product of the distance separating charges of equal magnitude and opposite sign, with the magnitude of the charge. The distance between the positive and negative centres called the bond length.

Thus, = μ = electric charge × bond length = q × d

As q is in the order of 10–10 esu and d is in the order of 10–8 cm, μ is the order of10–18 esu cm. Dipole moment is measured in ‘Debye’ unit (D)

1D = 10–18 es cm = 3.33 × 10–30 coulomb metre

Note:

  1. Generally as electronegativity difference increase in diatomic molecules, polarity of bond between the atoms increases therefore value of dipole moment increases.  

  2. Dipole moment is a vector quantity. 

  3. A symmetrical molecule is non-polar even though it contains polar bonds. For example, CO2, BF3, CCl4 etc. because summation of all bond moments present in the molecules cancel each other.

  4. Unsymmetrical non-linear polyatomic molecules have net value of dipole moment. For example, H2O, CH3OH, NH3 etc.?

Calculation of Resultant Bond Moments

1374_Polar bonds.JPG

 

Let AB and AC are two polar bonds inclined at an angle  θ their dipole moments are μ1 and μ2.

Resultant dipole moment may be calculated using vectorial method.

μ = √μ12 + μ22 + 2μ1μ2 cos θ

310_min. and max. resultant.JPG

when θ = 0 the resultant is maximum μR = μ1 + μ2
 

when, θ = 180°, the resultant is minimum μR = μ1 ∼ μ2

For example, CO2 has got dipole moment of zero. The structure of CO2 is.This is a highly symmetrical structure with a plane of symmetry passing through the carbon. The bond dipole of C–O is directed towards oxygen as it is the negative end. Here two equal dipoles acting in opposite direction cancel each other and therefore the dipole moment is zero. Similarly  dipole moment of CCl4 is zero while that of CHCl3 is non zero. . Explanation is again in geometry of the molecules both CCl4 & CHCl3 have tetrahedral structure but CCl4 is symmetrical while CHCl3 is non-symmetrical.

981_Symmetrical and Non-symmetrical.JPG

Due to the symmetrical structure of CCl4 the resultant of bond dipoles comes out to be zero. But in case of CHCl3 it is not possible as the presence of hydrogen introduces some dissymmetry.